Solve using the quadratic formula:

x2 + 4x +3=0


1
Expert's answer
2021-03-01T06:37:15-0500

x2 + 4x +3=0


let's use the discriminant formula

D = b2 - 4ac 


a=1

b=4

c=3


D = b2 - 4ac = 42-4*1*3=16-12=4


now let's find the x1 and x2 using the discriminant

x1= "(-b+\sqrt{\smash[b]{D}})/2a" = "(-4+\sqrt{\smash[b]{4}}))/2*1=(-4+2)/2=-2/2=-1"

x2= "(-b-\sqrt{\smash[b]{D}})/2a" ="(-4-\sqrt{\smash[b]{4}}))/2*1=(-4-2)/2=-6/2=-3"


The answer is

x1= -1

x2= -3